ECAL: problematic cells

Cells with (presumably) damaged optical fibers of LED Monitoring system

Week or absent LED signal.

Side Section X/Y R/O Crate/Slot/Channel Comment
C O 17/16 10/02/17  
18/17 10/02/26  
01/56 08/18/17  
31/57 10/20/31  
08/55 09/18/08  
A I 55/44 15/17/23  
32/14 15/02/00  
33/14 15/02/01  
32/16 15/02/16  
55/35 15/13/15  
55/36 15/13/23  
50/38 15/14/02  
53/32...I 55/34 15/12/(21..23;29..31 and 15/13/(5..7) LED A3.11.4...entire bundle is damaged
53/38...55/40 15/14/(5..7;13..15;21..23) LED A3.08.4, entire bundle is damaged
O 37/21 19/03/29  
36/54...39/57 19/14/(4..7; 12..15; 20..23; 28..31) LED A1.05.1...entire bundle is damaged

"Hot" cells

Abnormally high gain value, which can also vary in time. Cell should be masked on the level of trigger / readout systems except it is marked by (*)

Side Section X/Y R/O Crate/Slot/Channel Date Comment
C O 23/12 10/01/23 11/08/2008 Problem with PM HV control
A I 37/17(*)   26/08/2008 Problem with PM HV configuration. HV in DB is set to zero.
M 53/29 18/11/13 11/08/2008 Problem with PM HV control
54/51 18/18/30 11/08/2008 Problem with PM HV control

"Dead" cells

Malfunctioning C-W base, problem in readout FEB or absent contact in-between

Characterized by very narrow pedestal distribution or strange ADC spectrum

Side Section X/Y R/O Crate/Slot/Channel Date Comment
C O 29/21     Also no hits in "cosmic" data (run 30436)
O 15/37 09/11/31   Also no hits in "cosmic" data(run 30436)
I 28/14 13/02/04   Also no hits in "cosmic" data(run 30436)
O 36/10     ADC spectrum: Narrow pedestal + narrow peak around
48/22 20/07/00   Also no hits in "cosmic" data(run 30436)
48/34     Double LED peak; pedestal OK

Breakdown of capacitance in one of multiplication stages of C-W base.

Characterized by large pick-up noise and increased power consumption from Agilent PS (~3mA / defective base vs ~0.5 mA / operational base). Such base can from time to time produce large current pulse, which can affect trigger system. Therefore PM HV value for these cells has to be kept equal to zero.

Side Section X/Y Control box #/Address R/O Crate/Slot/Channel Date Comment
C O 01/43 01/054 08/13/09 15/08/2008
02/20 01/117 08/06/18 15/08/2008
09/34 03/063 09/11/01 28/07/2008
12/52 04/005 09/17/20 28/07/2008
16/08 08/033 10/00/16 28/07/2008
17/20 08/061 10/03/17 15/08/2008
30/45 14/172 10/17/30 28/07/2008
M 00/34 05/017 11/09/16 28/07/2008
00/51 05/000 11/14/24 28/07/2008
03/26 05/145 11/07/19 28/07/2008
07/36 06/095 11/11/07 28/07/2008
10/34 07/017 12/09/18 15/08/2008
11/49 07/042 12/14/11 28/07/2008
12/26 07/105 12/07/20 28/07/2008
15/24 15/027 12/07/07 15/08/2008
15/27 15/024 12/07/31 15/08/2008
20/49 15/124 12/18/12 28/07/2008
23/16 15/193 11/18/07 28/07/2008
26/48 16/043 12/20/02 28/07/2008
28/44 16/087 12/19/04 28/07/2008
30/48 16/123 12/20/06 28/07/2008
I 08/47 10/004 13/18/08 28/07/2008
09/18 10/075 13/08/01 28/07/2008
18/24 12/029 14/06/18 15/08/2008
20/45 12/086 14/12/28 26/08/2008
21/25 12/148 14/06/29 28/07/2008
22/30 12/183 14/08/06 28/07/2008
A I 32/17 20/198 15/02/24 15/08/2008
34/22 20/153 15/06/02 28/07/2008
35/25 20/130 15/06/27 28/07/2008
38/41 20/066 16/14/30 07/10/2008
42/20 21/153 16/03/18 28/07/2008
42/22 21/151 16/06/02 28/07/2008
42/32 21/139 16/08/18 28/07/2008
43/40 21/091 16/11/19 28/07/2008
44/14 21/079 16/02/04 28/07/2008
45/22 21/031 16/06/05 26/08/2008
45/39 21/012 16/11/13 07/10/2008
46/29 22/184 16/07/30 28/07/2008
48/28 22/105 15/11/16 28/07/2008
48/32 22/099 15/12/16 28/07/2008
52/44 23/127 15/17/20 28/07/2008
52/48 23/123 15/18/20 28/07/2008
M 38/15 17/034 17/00/30 28/07/2008
40/49 18/184 18/03/08 28/07/2008
49/41 26/170 18/14/09 28/07/2008
50/14 26/157 18/06/18 28/07/2008
53/39 26/012 18/13/29 28/07/2008
57/36 27/055 17/13/01 28/07/2008
58/36 27/015 17/13/02 28/07/2008
58/51 27/000 17/18/26 28/07/2008
59/36 28/175 17/13/03 08/08/2008
60/15 28/156 17/06/28 28/07/2008
61/33 28/098 21/09/29 26/08/2008
O 36/54 19/043 19/14/04 15/08/2008
37/55 19/002 19/14/13 15/08/2008
40/57 24/080 19/20/24 28/07/2008
46/18 25/063 19/09/06 28/07/2008
46/52 25/045 19/19/22 07/10/2008
46/21 25/060 19/09/30 28/07/2008
50/12 29/050 20/02/18 28/07/2008
51/48 29/009 20/14/19 28/07/2008
53/09 30/197 20/01/29 07/10/2008
53/26 30/091 20/08/05 15/08/2008
54/18 30/079 20/06/06 28/07/2008
54/52 30/045 20/17/22 28/07/2008
57/30 31/107 21/09/01 28/07/2008
63/53 32/004 21/17/31 28/07/2008
63/45 32/012 21/13/31 28/07/2008

Other

Absent LED, but C-W noise is OK (i.e. no problem for trigger)

Side Section X/Y R/O Crate/Slot/Channel Date Comment
C M 08/24     no hits in both LED/"cosmic" data (run 30436)
I 22/39  
A M 45/50     no hits in both LED/"cosmic" data(run 30436)
60/40     no hits in both LED / "cosmic" data(run 30436)
O 37/21     no hits in "cosmic" data(run 30436)

-- IrinaMachikhiliyan - 18 Aug 2008

Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r8 - 2008-10-08 - IrinaMachikhiliyan
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback